OpenStudy (anonymous):

Are these correct? A candle has the mass of 10 g and a volume of 16 cm3. What is it’s density(g/cm3)? .16g/cm3 A ball has the mass of 3 g and a volume of 9 cm3. What is it’s density (g/cm3) 3g/cm3 A computer has a mass of 20 g and a volume of 40 cm3. What is it’s density (g/cm3) 2g/cm3 What is the volume of a box with the density of 250 g/cm3 and a mass of 30 g? 8.3cm3 If any wrong I need help find how to do it

OpenStudy (chmvijay):

Density is =Mass/Volume

OpenStudy (chmvijay):

all are wrong

OpenStudy (chmvijay):

in most of cases you have done volume/mass please correct it

OpenStudy (anonymous):

if you do mass over volume you get what I got?

OpenStudy (chmvijay):

10/16=0.625 3/9=0.33 20/40=0.5 30/250=0.12 Remember always density values will be less than one for most of organic solvent some are exceptions are there